# Holistic Optimization Framework for FPGA Accelerators

### v3

@@ 8 F[i1][j1] += C[i][k] * D[k][j1];}
Listing 3: Pseudo Code of the task_3
Data-tiling and Unroll Factor  The intra-tile transformation, as explained in Section III, can divide either the original loop trip count or the original trip count with padding, thereby increasing the range of possibilities. Equation 1 ensures that the trip count of the intra-tile is a divisor of one of these two possibilities. The user has the option to constrain the padding using Equation 2, which simplifies the solution space for the NLP solver. For instance, in Listing 3, for the array $F$, the loop $j$ (line 7 in Listing 1) has been split into $j0$ (line 14 in Listing 2) and $j1$ (intra-tile). The trip count of the intra-tile loop $j1$, denoted as $TC^{j1}_{\text{intra}}=32$, does not evenly divide the original trip count $TC^{j}_{\text{ori}}=210$, but it does divide the trip count of the padded loop $TC^{j}=224$.
$\forall l\in\mathcal{L},TC^{l}_{intra}\%TC^{l}_{ori}==0||TC^{l}_{intra}\%TC^{l}==0$ (1)
$(opt)\forall l\in\mathcal{L},\exists n\in\mathbb{N}\leq N\in\mathbb{N},\text{ s.t. }TC^{l}=TC^{l}_{ori}+n$ (2)
Bit Width  $\mathcal{B}$ represents the possible bit width for the data type. Hence, if we have a bit width under 512 bits for float the set is $\{1,2,4,8,16\}$. Equation 3 computes the bit width for each array based on the last dimension of the data-tile transferred on-chip. For example, the array D in Fused task 1 is transferred in line 15 with a size of $222\times 32$, so $S^{last}_{D}=32$, which is divisible by 16. Therefore, it has a bit width of 16.
$\forall a\in\mathcal{A},\forall l\in\mathcal{L}_{a}^{last},\max_{b\in\mathcal{B}}BW_{a}=b\text{ s.t. }S^{last}_{a}\%b=0$ (3)
Permutation  Equation 4 ensures that the NLP selects a unique permutation for each loop that iterates through every statement. Equation 5 requires the NLP to choose identical permutations for loops that are shared by statements fused within the same task. For example, in Fused task 1, the loops iterating statement S2 can be permuted as ($i0$, $j0$) or ($j0$, $i0$), and similarly, loops iterating statement S3 can be permuted as ($i0$, $j0$) or ($j0$, $i0$). However, since the loops $i0$ and $j0$ iterate both S2 and S3, they must use the same permutation; either ($i0$, $j0$) for both or ($j0$, $i0$) for both.
$\forall s\in\mathcal{S},\forall i\in\mathcal{P}_{s},\sum_{p}p_{i}^{l}=1$ (4)
@@
$\forall(i_{0},i_{1})\in\mathcal{P}_{ft_{0}}\times\mathcal{P}_{ft_{1}},\forall l\in\mathcal{L},p_{i0}^{l}=p_{i1}^{l}$ (5)
Transfer and Reuse  Equation 6 permits the selection of a single level where each array can be defined (and reused) and transferred. Equation 7 constrains that the definition of the array must occur lexicographically before or at the same time as the transfer.
The array $E$, defined on line 24 in Listing 2, is defined before any loops, so $d_{E,0}=1$ (0 indicating it is defined before any loops). However, it is transferred under the loop $i0$, so $t_{E,i0}=1$. Equation 7 simply means that the definition of array E should occur before or at the same level as the transfer. We cannot transfer E under loop $i0$ if E is defined under $k0$. Similarly, in Listing 2, the array $A$ is defined and transferred in line 4, with $d_{A,i0}=1$ and $t_{A,i0}=1$.
$\forall a\in\mathcal{A},\sum l\in\mathcal{L}_{inter},t_{a,l}=1,\sum l\in\mathcal{L}_{inter},d_{a,l}=1$ (6)
$\forall a\in\mathcal{A},\sum(l_{0},l_{1})\in\mathcal{L}_{inter}^{2}|d_{a,l_{0}},t_{a,l_{1}}=1,\text{ then }l_{0}\preccurlyeq l_{1}$ (7)
On-chip Memory  Equation 8 constrains the footprint of the array to be within the available resources, based on where the array is defined, the number of double buffers used and the footprint of the array transferred at this level.
$\sum_{a\in\mathcal{A}}\sum_{l\in\mathcal{L}}d_{a,l}\times f_{a,l}\times N_{a}\leq Mem,$ (8)
@@
Array Partitioning  Equation 9 limits the maximum partitioning of each array. This partitioning is crucial as it impacts the maximum unroll factor, necessitating the distribution of data across different BRAM banks under fully unrolled loops, thereby influencing the utilization of BRAMs. Equation 10 computes the array partitioning needed for each array based on the trip count of the fully unrolled intra-tile loops.
Array D in Listing 3 is traversed by two unrolled loops: $k1$, which iterates 3 times ($AP_{D,0}=3$), and $j1$, which iterates 32 times ($AP_{D,1}=32$). Therefore, the total number of partitions needed is $3\times 32=96$. Consequently, these 96 values of F are stored in different banks, allowing all of them to be accessed in parallel. However, this value must be less than or equal to $max_{part}$.
$\forall a\in\mathcal{A},\prod_{d\in\mathbb{N}}AP_{a,d}\leq max_{part}$ (9)
$\forall a\in\mathcal{A},\forall d\in\mathbb{N},\forall l\in C_{a_{d}},AP_{a,d}=TC^{l}_{intra}==0$ (10)
DSP Utilization  Equation 11 constrains the number of DSPs used based on the available DSP resources. In contrast to [59, 58, 60], we utilize pessimistic DSP utilization. This approach is necessary because concurrent execution and resource reuse between tasks are not feasible when two tasks can run simultaneously.
Given $DSP_{+}=2$, $DSP_{*}=3$, and $II_{S3}=3$, the DSP usage for Task 3 is calculated as $(2+3)\times 1824$, accounting for the unroll factor. However, since the loop is pipelined with $II=3$, the HLS compiler optimizes resource usage, effectively reducing the DSP count by approximately dividing by $II$.
$\sum_{op\in\{+,-,*,/\}}\sum_{s\in\mathcal{S}}(DSP_{s_{op}}/II_{s})\times\prod_{l\in\mathcal{L_{s}}}TC^{l}_{intra}\leq DSP$ (11)
SLR Selection  For each task, the NLP determines the SLR on which the task will be implemented (Equation 12). Additionally, Equations 8 and 11 are applied to each SLR to manage resource allocation per SLR effectively.
$\forall{t\in\mathcal{T}},slr_{t}\in\llbracket 0,SLR\llbracket$ (12)
IV-B Objective Function
@@
T0T1T2T3T4T5
Figure 1: Dataflow graph of 3mm maximally distributed
The latency of the intra-task for a statement $s$ is determine similarly to [59, 58, 60] with Equation 13. And if the intra-task is iterated by the reduction loops in the inter-tile the latency is represented by Equation 14.
$Lat_{intra}=IL_{par}+IL_{seq}\times\log_{2}(\prod_{l\in\mathcal{L}^{red}_{s}}TC^{l}_{intra})$ (13)
$Lat=Lat_{intra}+II\times(\prod_{l\in\mathcal{L}^{red}_{s}}TC^{l}_{inter}-1)$ (14)
@@
$Lat_{n}=\max(Lat_{n+1},\frac{f_{a,l}}{BS_{a}})+Lat_{n+1}+\frac{f_{a,l}}{BS_{a}}\times\alpha$ (15)
@@ V Code Generation
Prometheus takes as input affine C/C++ code and automatically produces an HLS-C++ file, OpenCL host files, and all necessary files for code verification, RTL simulation, and bitstream generation. The NLP described in Section IV gives the parameters of the space such as loop order, tiling factor, etc. However, in order to be efficient, the code generation needs some specific rules.
@@
Communication with Off-Chip Memory  To enable overlapping communication and computation, a load function transfers data on-chip using FIFOs, ensuring effective overlap management by the compiler. Furthermore, to guarantee that the load operation transfers data with the correct bit width, we automatically restructure the data in off-chip memory to enable sequential loading. Once data accumulates in the FIFO, a read function moves it to the shared data-tile buffer, whose size is determined by the NLP. For outputs, a write function transfers data from the buffer to a FIFO, and a store function writes it back to off-chip memory.
@@
Communication between the Fused Task  The communication within the same fused task is not required as they use shared buffers. Similarly, for communication with off-chip memory, we choose to use FIFOs to facilitate communication between fused tasks, thereby simplifying the overlap.
Intra-Task  Each intra-task, corresponding to the intra-tile selected by the NLP, is implemented as an independent, fully unrolled function with no communication with off-chip memory. Data for these tasks resides on-chip. If the task includes reduction loops, inter-tile reduction loops are incorporated and pipelined within the function. The pipeline’s II is strictly greater than 1 due to reduction loop dependencies, but other operations within the statements are effectively pipelined. Padding is managed at the intra-tile level. For non-reduction loops, padding does not alter the intra-task, allowing computation of padding values without unnecessary resource usage for minor latency improvements. In contrast, for reduction loops, full tile computations are performed first, followed by adjusting the intra-tile loop to handle padding and compute partial tiles accurately.
@@
Inter-Tile Loop  For each fused task and each inter-tile loop, we generate independent functions to facilitate efficient double or triple buffering. Using information from the NLP, we determine which arrays are defined and which are transferred at each level of granularity. Once we identify the data being transferred, we implement double buffering if only reads are involved, or triple buffering if both reads and writes are required. This strategy allows us to overlap the operations of reading, storing, and computing at the innermost level, optimizing both communication and computation overlap.
@@
Concurrent Execution  With pragma dataflow, independent tasks run concurrently without rewriting. Computation begins as soon as the receiving task’s shared buffer has enough data.
@@
SLR  The NLP determines the SLR on which each task will be executed. A separate C++ file is generated for each SLR, and data transfers between SLRs are handled using ap_axiu.
VI Evaluation
VI-A Setup and Experimental Evaluation
@@
We evaluated our method using kernels from Polybench/C 4.2.1 [56] with medium-sized datasets and single-precision floating-point computations. The selected kernels represent both memory-intensive and computation-intensive scenarios. Medium problem sizes were chosen to balance demonstration of efficacy and the feasibility of time-consuming RTL evaluations. NLP problems were solved using the AMPL description language and the Gurobi solver (version 11.0.0) with the qp:nonconvex=2 option for non-convex quadratic objectives and constraints. Evaluations included RTL simulation and on-board execution using the Alveo U55C FPGA, with a targeted frequency of 220 MHz. RTL simulation provided accurate latency estimates, contrasting with the overly optimistic Vitis HLS reports that assume perfect task overlapping with dataflow pragma. The generated code from the frameworks are compiled using AMD/Xilinx Vitis HLS 2023.2 using the Vitis flow [4]. This flow assumes that data initially resides off-chip and has a default latency of 64 cycles to bring onto on-chip memory. All frameworks utilize ”unsafe math” optimizations, enabling commutative/associative reduction operations at the expense of precision.
@@
The objective of this evaluation is to demonstrate the capability of our framework to generate code with high QoR, whether for memory-bound or computation-bound kernels. To achieve this, we comare our work with Allo, ScaleHLS, Sisyphus, and AutoDSE. We execute the designs for these frameworks using identical parameters as outlined in Section VI-A. For ScaleHLS [79] and Allo [14], we apply the same method as our framework by padding memory transfers to attain the maximal bit width (512 bits) from off-chip to on-chip memory. Their kernels assume data are already in on-chip memory, so we modified their code to include off-chip to on-chip data transfers. Allo’s 2mm and 3mm kernels already handle this transfer, requiring no modifications. We left Sisyphus and AutoDSE unchanged, as they already optimize bit width according to the problem size. We use AutoDSE with the bottleneck method, setting a DSE timeout of 1,000 minutes and an HLS synthesis timeout of 180 minutes per task. For Allo-generated designs, we utilize kernels from their artifact repository since Allo does not employ a DSE [1]. Sisyphus generated designs are generated using parameters consistent with the paper [60], and ScaleHLS designs are compiled using their open-source compiler [2]. For RTL simulation, we assume that the frameworks can utilize all the resources on the U55C FPGA with a constraint of partitioning any array of 1,024 due to AMD/Xilinx limitations. For on-board FPGA evaluations, we consider two scenarios. The first scenario utilizes 60% of one Super Logic Region (SLR), equivalent to 20% of total board resources, for all frameworks. The second scenario is unique to our framework, leveraging all three SLRs, with 60% utilization per SLR. Most frameworks are not place-and-route aware, leading to congestion issues that prevent bitstream generation. Limiting frameworks to one SLR increases the likelihood of meeting timing requirements. AutoDSE, for instance, lacks dataflow support and applies pragmas within a single function, making multi-SLR bitstream generation impossible without manual intervention to split the function across SLRs. If congestion persists within an SLR, we adjust the NLP constraints to address the issue and regenerate the HLS-C++ file in just a few minutes.
VI-B Comparison
The average time to solve the NLP and generate the designs presented in this section is 31.7 seconds, with a maximum of 178.416 seconds for 3mm with 3 SLRs.

### v4

@@ 8 F[i1][j1] += C[i][k] * D[k][j1];}
Listing 5: Implementation of Task 3 in the 3mm Kernel (from Listing 4). This listing illustrates the structure of Task 3, where the intra-tile computation is fully unrolled to expose fine-grained parallelism. The unrolling enables concurrent execution of loop iterations.
4.1.1. Data-tiling and Unroll Factor
The intra-tile transformation, as explained in Section 3, can divide either the original loop trip count or the original trip count with padding, thereby increasing the range of possibilities. Equation 1 ensures that the trip count of the intra-tile is a divisor of one of these two possibilities. The user has the option to constrain the padding using Equation 2, which simplifies the solution space for the NLP solver. For instance, in Listing 5, for the array $F$, the loop $j$ (line 7 in Listing 3) has been split into $j0$ (line 14 in Listing 4) and $j1$ (intra-tile). The trip count of the intra-tile loop $j1$, denoted as $TC^{j1}_{\text{intra}}=32$, does not evenly divide the original trip count $TC^{j}_{\text{ori}}=210$, but it does divide the trip count of the padded loop $TC^{j}=224$.
(1) $\forall l\in\mathcal{L},TC^{l}_{intra}\%TC^{l}_{ori}==0||TC^{l}_{intra}\%TC^{l}==0$
(2) $(opt)\forall l\in\mathcal{L},\exists n\in\mathbb{N}\leq N\in\mathbb{N},\text{ s.t. }TC^{l}=TC^{l}_{ori}+n$
4.1.2. Bit Width
$\mathcal{B}$ denotes the number of elements that can be transferred simultaneously, determined by the bit width and the data type. Hence, if we have a bit width under 512 bits for float the set is $\{1,2,4,8,16\}$. Equation 3 computes the bit width for each array based on the last dimension of the data-tile transferred on-chip. For example, the array D in Fused task 1 is transferred in line 15 with a size of $222\times 32$, so $S^{last}_{D}=32$, which is divisible by 16. Therefore, it has a bit width of 16.
(3) $\forall a\in\mathcal{A},\forall l\in\mathcal{L}_{a}^{last},\max_{b\in\mathcal{B}}BW_{a}=b\text{ s.t. }S^{last}_{a}\%b=0$
4.1.3. Permutation
Equation 4 requires the NLP to choose identical permutations for loops that are shared by statements fused within the same task. For example, in Fused task 1, the loops iterating statement S2 can be permuted as ($i0$, $j0$) or ($j0$, $i0$), and similarly, loops iterating statement S3 can be permuted as ($i0$, $j0$) or ($j0$, $i0$). However, since the loops $i0$ and $j0$ iterate both S2 and S3, they must use the same permutation; either ($i0$, $j0$) for both or ($j0$, $i0$) for both.
@@ 4.1.3. Permutation
(4) $\forall(i_{0},i_{1})\in\mathcal{P}_{ft_{0}}\times\mathcal{P}_{ft_{1}},\forall l\in\mathcal{L},p_{i0}^{l}=p_{i1}^{l}$
4.1.4. Transfer and Reuse
Equation 5 permits the selection of a single level where each array can be defined (and reused) and transferred. Equation 6 constrains that the definition of the array must occur lexicographically before or at the same time as the transfer.
The array $E$, defined on line 24 in Listing 4, is defined before any loops, so $d_{E,0}=1$ (0 indicating it is defined before any loops). However, it is transferred under the loop $i0$, so $t_{E,i0}=1$. Equation 6 simply means that the definition of array E should occur before or at the same level as the transfer. We cannot transfer E under loop $i0$ if E is defined under $k0$. Similarly, in Listing 4, the array $A$ is defined and transferred in line 4, with $d_{A,i0}=1$ and $t_{A,i0}=1$.
(5) $\forall a\in\mathcal{A},\sum l\in\mathcal{L}_{inter},t_{a,l}=1,\sum l\in\mathcal{L}_{inter},d_{a,l}=1$
(6) $\forall a\in\mathcal{A},\sum(l_{0},l_{1})\in\mathcal{L}_{inter}^{2}|d_{a,l_{0}},t_{a,l_{1}}=1,\text{ then }l_{0}\preccurlyeq l_{1}$
4.1.5. On-chip Memory
Equation 7 constrains the footprint of the array to be within the available resources, based on where the array is defined, the number of double buffers used and the footprint of the array transferred at this level.
(7) $\sum_{a\in\mathcal{A}}\sum_{l\in\mathcal{L}}d_{a,l}\times f_{a,l}\times N_{a}\leq Mem,$
@@ 4.1.5. On-chip Memory
4.1.6. Array Partitioning
Equation 8 limits the maximum partitioning of each array. This partitioning is crucial as it impacts the maximum unroll factor, necessitating the distribution of data across different BRAM banks under fully unrolled loops, thereby influencing the utilization of BRAMs. Equation 9 computes the array partitioning needed for each array based on the trip count of the fully unrolled intra-tile loops.
Array D in Listing 5 is traversed by two unrolled loops: $k1$, which iterates 3 times ($AP_{D,0}=3$), and $j1$, which iterates 32 times ($AP_{D,1}=32$). Therefore, the total number of partitions needed is $3\times 32=96$. Consequently, these 96 values of F are stored in different banks, allowing all of them to be accessed in parallel. However, this value must be less than or equal to $max_{part}$.
(8) $\forall a\in\mathcal{A},\prod_{d\in\mathbb{N}}AP_{a,d}\leq max_{part}$
(9) $\forall a\in\mathcal{A},\forall d\in\mathbb{N},\forall l\in C_{a_{d}},AP_{a,d}=TC^{l}_{intra}==0$
4.1.7. DSP Utilization
Equation 10 constrains the number of DSPs used based on the available DSP resources. In contrast to (Pouget et al., 2025a, 2024, b), we utilize pessimistic DSP utilization. This approach is necessary because concurrent execution and resource reuse between tasks are not feasible when two tasks can run simultaneously. Given $DSP_{+}=2$, $DSP_{*}=3$, and $II_{S3}=3$, the DSP usage for Task 3 is calculated as $(2+3)\times 1824$, accounting for the unroll factor. However, since the loop is pipelined with $II=3$, the HLS compiler optimizes resource usage, effectively reducing the DSP count by approximately dividing by $II$.
(10) $\sum_{op\in\{+,-,*,/\}}\sum_{s\in\mathcal{S}}(DSP_{s_{op}}/II_{s})\times\prod_{l\in\mathcal{L_{s}}}TC^{l}_{intra}\leq DSP$
4.1.8. SLR Selection
For each task, the NLP determines the SLR on which the task will be implemented (Equation 11). Additionally, Equations 7 and 10 are applied to each SLR to manage resource allocation per SLR effectively.
(11) $\forall{t\in\mathcal{T}},slr_{t}\in\llbracket 0,SLR\llbracket$
4.2. Objective Function
@@
PrometheusT0T1T2T3T4T5
Figure 2. Maximally distributed dataflow graph of the 3mm kernel, where each node represents a task and each edge indicates a data dependency between tasks.
The latency of the intra-task for a statement $s$ is determine similarly to (Pouget et al., 2025a, 2024, b) with Equation 12. And if the intra-task is iterated by the reduction loops in the inter-tile the latency is represented by Equation 13.
(12) $Lat_{intra}=IL_{par}+IL_{seq}\times\log_{2}(\prod_{l\in\mathcal{L}^{red}_{s}}TC^{l}_{intra})$
(13) $Lat=Lat_{intra}+II\times(\prod_{l\in\mathcal{L}^{red}_{s}}TC^{l}_{inter}-1)$
@@
(14) $Lat_{n}=\max(Lat_{n+1},\frac{f_{a,l}}{BS_{a}})+Lat_{n+1}+\frac{f_{a,l}}{BS_{a}}\times\alpha$
@@ 5. Code Generation
Prometheus takes as input affine C/C++ code and automatically produces an HLS-C++ file, OpenCL host files, and all necessary files for code verification, RTL simulation, and bitstream generation. The NLP described in Section 4 gives the parameters of the space such as loop order, tiling factor, etc. However, in order to be efficient, the code generation needs some specific rules.
5.1. Communication with Off-Chip Memory
@@
To enable overlapping communication and computation, a load function transfers data on-chip using FIFOs, ensuring effective overlap management by the compiler. Furthermore, to guarantee that the load operation transfers data with the correct bit width, we automatically restructure the data in off-chip memory to enable sequential loading. Once data accumulates in the FIFO, a read function moves it to the shared data-tile buffer, whose size is determined by the NLP. For outputs, a write function transfers data from the buffer to a FIFO, and a store function writes it back to off-chip memory.
5.2. Communication between the Fused Task
@@
The communication within the same fused task is not required as they use shared buffers. Similarly, for communication with off-chip memory, we choose to use FIFOs to facilitate communication between fused tasks, thereby simplifying the overlap.
5.3. Intra-Task
Each intra-task, corresponding to an intra-tile selected by the NLP, is implemented as a fully unrolled, independent function without communication with off-chip memory. Data for these tasks resides entirely on-chip.
If the task involves reduction loops, inter-tile reduction loops are integrated into the function and pipelined. While the pipeline initiation interval (II) is greater than 1 due to reduction dependencies, other operations in the statement are pipelined efficiently.
Padding is handled at the intra-tile level. Non-reduction loops remain unchanged, allowing computation of padding values without excessive resource usage. For reduction loops, full tiles are computed first, and the intra-tile loop is adjusted to handle padding for partial tiles accurately.
5.4. Inter-Tile Loop
@@
For each fused task and each inter-tile loop, we generate independent functions to facilitate efficient double or triple buffering. Using information from the NLP, we determine which arrays are defined and which are transferred at each level of granularity. Once we identify the data being transferred, we implement double buffering if only reads or writes are involved, or triple buffering if both reads and writes are required. This strategy allows us to overlap the operations of reading, storing, and computing at the innermost level, optimizing both communication and computation overlap.
5.5. Concurrent Execution
@@
Using the dataflow pragma, independent tasks execute concurrently without requiring manual rewrites. Computation in a receiving task begins as soon as its shared buffer contains sufficient data.
5.6. SLR Management
@@
The NLP determines the SLR ID for each task, specifying where it will be executed. Prometheus generates a separate C++ file for each SLR, and data transfers between SLRs are managed via ap_axiu streams, ensuring efficient communication and minimizing transfer overhead.
5.7. Design Regeneration
In cases where bitstream generation fails for a given design, we support automatic regeneration of the design. Several strategies can be applied, such as tightening resource constraints or reducing the maximum unrolling factor. By reducing the available resources, the design becomes smaller, which can help alleviate congestion. Thanks to the flexibility of our NLP-based approach, we can retain parts of the previous solution—such as the SLR assignment—and selectively restrict resources only for the specific task or group of tasks responsible for the congestion. This is the method we employ to regenerate designs when bitstream generation fails.
6. Evaluation
6.1. Setup
@@
We evaluated our method using kernels from Polybench/C 4.2.1 (Pouchet and Yuki, 2024) with medium-sized datasets and single-precision floating-point computations. The selected kernels represent both memory-intensive and computation-intensive scenarios. Medium problem sizes were chosen to balance demonstration of efficacy and the feasibility of time-consuming RTL evaluations. Due to Allo’s lack of automatic code generation, we limited our experiments to the subset of PolyBench kernels provided in its artifact evaluation package (all, 2024), as results for other kernels were not available. NLP problems were solved using the AMPL description language and the Gurobi solver (version 11.0.0) with the qp:nonconvex=2 option for non-convex quadratic objectives and constraints. Evaluations included RTL simulation and on-board execution using the Alveo U55C FPGA, with a targeted frequency of 220 MHz. RTL simulation provided accurate latency estimates, contrasting with the overly optimistic Vitis HLS reports that assume perfect task overlapping with dataflow pragma. The generated code from the frameworks are compiled using AMD/Xilinx Vitis HLS 2023.2 using the Vitis flow (AMD/Xilinx, 2024). This flow assumes that data initially resides off-chip and has a default latency of 64 cycles to bring onto on-chip memory. All frameworks utilize ”unsafe math” optimizations, enabling commutative/associative reduction operations at the expense of precision.
6.2. Experimental Evaluation
The objective of this evaluation is to demonstrate the capability of our framework to generate code with high QoR, whether for memory-bound or computation-bound kernels. To achieve this, we compare our work with Allo, ScaleHLS, Sisyphus, AutoDSE and Stream-HLS.
For ScaleHLS (Ye et al., 2022), Allo (Chen et al., 2024), and Stream-HLS (Basalama and Cong, 2025), their kernels assume that data is already present in on-chip memory. To ensure a fair comparison, we modified their code to incorporate off-chip to on-chip data transfers. However, Allo’s 2mm and 3mm kernels already include this transfer mechanism, requiring no modifications. We left Sisyphus and AutoDSE unchanged, as they already optimize bit width according to the problem size. We use AutoDSE with the bottleneck method, setting a DSE timeout of 1,000 minutes and an HLS synthesis timeout of 180 minutes per task. For Allo-generated designs, we utilize kernels from their artifact repository since Allo does not employ a DSE (all, 2024). Sisyphus generated designs are generated using parameters consistent with the paper (Pouget et al., 2025b). We observe some differences in Sisyphus results, as our evaluation is conducted on a different board. The design variations between boards may lead to differences in performance.
@@ 6.2. Experimental Evaluation
For RTL simulation, we assume that the frameworks can utilize all the resources on the U55C FPGA with a constraint of partitioning any array of 1,024 due to AMD/Xilinx limitations. For on-board FPGA evaluations, we consider two scenarios. The first scenario utilizes 60% of one Super Logic Region (SLR), equivalent to 20% of total board resources, for all frameworks. The second scenario is unique to our framework, leveraging all three SLRs, with 60% utilization per SLR. Most frameworks are not place-and-route aware, leading to congestion issues that prevent bitstream generation. Limiting frameworks to one SLR increases the likelihood of meeting timing requirements. AutoDSE, for instance, lacks dataflow support and applies pragmas within a single function, making multi-SLR bitstream generation impossible without manual intervention to split the function across SLRs. If congestion persists within an SLR, we adjust the NLP constraints to address the issue and regenerate the HLS-C++ file in just a few minutes.
It is worth noting that there is a difference in results between RTL simulation and on-board evaluation. This difference is expected, as RTL simulation involves fewer constraints, allowing for more aggressive unrolling and full utilization of available resources. In contrast, on-board evaluation is more complex, and excessive unrolling can easily lead to congestion. Therefore, we compare RTL simulation results with those of other frameworks to demonstrate our performance improvements. More importantly, we show that our approach generates code that can actually be implemented on board.
6.3. Comparison